OpenStudy (anonymous):

How can I get 2(n+1)=2^(n+1) - 2 to equal the same thing on both sides?

OpenStudy (anonymous):

whats the ^ on the 2nd part?

OpenStudy (anonymous):

\[2(n+1)=2^{n+1}-2\]

OpenStudy (anonymous):

2(n+1)+2=2^(n+1) 2n+4=2^(n+1) log(2n+4)=(n+1)log2 (log(2n+4))/(n+1)=log2 log((2n+4)^(1/(n+1))=log2 ((2n+4)^(1/(n+1))=2